THUMBNAIL(1) THUMBNAIL(1) NAME thumbnail−createaTIFFfilewiththumbnailimages SYNOPSIS thumbnail[options]input.tifoutput.tif DESCRIPTION thumbnailisaprogramwrittentoshowhowonemightusetheSubIFDtag(#330)tostore thumbnailimages.thumbnailcopiesaTIFFClassFfacsimilefiletotheoutputfileandforeach imagean8-bitgreyscalethumbnailsketch.Theoutputfilecontainsthethumbnailimagewiththe associatedfull-resolutionpagelinkedbelowwiththeSubIFDtag. Bydefault,thumbnailimagesare216pixelswideby274pixelshigh.Pixelsarecalculatedby samplingandfilteringtheinputimagewitheachpixelvaluepassedthroughacontrastcurve. OPTIONS −wSpecifythewidthofthumbnailimagesinpixels. −hSpecifytheheightofthumbnailimagesinpixels. −cSpecifyacontrastcurvetoapplyingeneratingthethumbnailimages.Bydefaultpixels valuesarepassedthroughalinearcontrastcurvethatsimplymapsthepixelvalueranges. Alternativecurvesare:exp50fora50%exponentialcurve,exp60fora60%exponential curve,exp70fora70%exponentialcurve,exp80fora80%exponentialcurve,exp90for a90%exponentialcurve,expforapureexponentialcurve,linearforalinearcurve. BUGS Therearenooptionstocontroltheformatofthesavedthumbnailimages. SEEALSO tiffdump(1),tiffgt(1),tiffinfo(1),libtiff(3) September26,1994 1